### v3

# Multi-Agent Target Assignment and Path Finding for Intelligent Warehouse: A Cooperative Multi-Agent Deep Reinforcement Learning Perspective

Abstract
Multi-agent target assignment and path planning (TAPF) are two key problems in intelligent warehouses. However, most literature only addresses one of these two problems separately. This paper proposes a method to simultaneously solve target assignment and path planning from the cooperative multi-agent deep reinforcement learning (RL) perspective. This paper models the TAPF problem for intelligent warehouses to cooperative multi-agent deep RL and simultaneously addresses task assignment and path finding based on multi-agent deep RL. Furthermore, previous literature rarely considers the physical dynamics of agents. In this paper, the physical dynamics of the agents are considered. This is more in line with actual robot movement. Experimental results show that the proposed method performs well in various task settings, which means that the target assignment is solved reasonably well and the planned path is almost the shortest. Moreover, the proposed method is more time-efficient than baselines.
I Introduction
With the rapid development of logistics-related industries, new opportunities, and challenges have been proposed for intelligent warehousing [1, 2]. Traditional warehouse technologies use conveyor belts and other equipment to complete material handling. They are inflexible and not easily extensible. Intelligent warehousing utilizes a multi-agent system to deliver goods to designated locations, greatly improving efficiency. The intelligent warehouse system is divided mainly into the order fulfillment system, such as the kiva system [3] and the logistics sorting center [4]. In order fulfillment systems, mobile agents move inventory pods to inventory stations. Then, workers take goods from inventory pods, and agents move the inventory pods to their original locations. In logistics sorting centers, agents take goods from loading stations and deliver them to chutes in the center of the warehouse. The approach proposed in this study is based on the logistics sorting center.
Task assignment and path finding (TAPF) are two important processes in intelligent warehousing. The system first assigns a specified task to each agent according to the order requirements. Then, the agent transports the goods from the origin to the destination and ensures that the path does not conflict with other agents. The task assignment and path planning problem is typically NP-hard [5], which has a large search space. Thus, directly solving this problem is difficult. In general, two steps are required to solve the TAPF problem. The first step is multi-agent task assignment (MATA), assigning tasks to agents without considering possible path conflicts between agents. The second step is path planning for all agents by using multi-agent path finding (MAPF) algorithms. Numerous studies in the literature study MATA and MAPF; we describe these methods in detail in the next section. Although solving TAPF separately reduces the difficulty of the total problem, this type of method ignores the mutual influence between task assignment and path planning. Reasonable task assignments can reduce warehouse agents’ path length, improve operation efficiency, and help avoid path conflicts between agents. Therefore, it is necessary to solve the TAPF problems together. In the TAPF problem, we assume that each idle agent and inventory pod are homogeneous. Thus, we can assign any task to any agent.
Recently, deep reinforcement learning (RL) has received great attention due to its remarkable successes in widely challenging domains, such as robotic control [6, 7, 8], multi-agent problems [9], and video games [10]. In this study, we introduce cooperative multi-agent deep RL to solve the TAPF problem and address it simultaneously.
The main contributions of this paper can be summarized as follows:
  • This paper proposes a method to simultaneously solve target assignment and path planning from the cooperative multi-agent deep RL perspective. This paper models the TAPF problem for intelligent warehouses to cooperative multi-agent deep RL and simultaneously solves task assignment and path finding based on the multi-agent deep RL algorithm.
  • Furthermore, in this paper, the physical dynamics of the agents are considered in the path planning phase. This is more in line with actual robot movement.
  • Experimental results show that the proposed method performs well in various task settings. This means the target assignment is solved reasonably, and the planned path is almost the shortest. Furthermore, the proposed method is more time-efficient than the baselines.
II Related Work
II-A Multi-Agent Task Assignment
Through multi-agent task assignment algorithms, we can maximize the utilization of warehouse resources [11]. Currently, multi-agent task assignment algorithms in intelligent warehouses can be classed into centralized and distributed classes according to the management mode. In the centralized task assignment class, a central control system is set up, which is responsible for task assignment, assigning tasks assigned to agents for execution [12]. Classical centralized task assignment algorithms include: Hungarian algorithm [13], tabu search algorithm [14], genetic algorithm [12], etc. In the distributed assignment class, each agent in the warehouse plans its own task sequence according to tasks and environmental information [15]. This class method effectively reduces the load of the central control system and is more flexible and adaptable, but it may not find the global optimal solution [16]. Furthermore, the distributed assignment class mainly includes learning-based methods [17] and market auction methods [18].
II-B Multi-Agent Path Finding
In recent years, MAPF has become a hot research direction in computer science and robotics. Classical MAPF algorithms can be divided into optimal and sub-optimal types according to whether the solution results meet the optimality. The optimal typical MAPF algorithms include A*-based [19], conflict-based search [20], increasing cost tree search based [21], and compilation-based [22] methods. Sub-optimal classical MAPF algorithms include search-based [23], rule-based [24], compilation-based [25], bounded conflict-based search [26], and genetic algorithm-based [27] methods. However, traditional MAPF algorithms have poor real-time performance. Therefore, numerous researchers are beginning to study MAPF based on deep RL [28] to solve this problem. PRIMAL [29] is a typical MAPF algorithm based on deep RL. However, PRIMAL still solves the multi-agent problem by single-agent deep RL method [29]. In this study, we propose a method to simultaneously address target assignment and path planning from the cooperative multi-agent deep RL perspective. Furthermore, in the simulation environment, the above-mentioned various MAPF algorithms rarely consider the physical dynamics of agents. In this study, the physical dynamics factors of the agent are considered.
II-C Cooperative Multi-Agent Deep RL
Cooperative multi-agent RL (MARL) deals with systems consisting of several agents that interact in a common environment. Cooperative multi-agent RL aims to learn a policy to make all agents work together to complete a common goal. Multi-agent deep RL has received great attention due to its capability to allow agents to learn to make decisions in multi-agent environments through interactions. Independent Q-learning (IQL) [30] solves the multi-agent RL problem by decomposing it into a collection of simultaneous single-agent RL problems that share the same environment. However, IQL cannot solve the nonstationarity [31] problem caused by the changing policies of other agents. Value-Decomposition Networks (VDN) [32] proposes to learn to linearly decompose a team value function into a series of per-agent value functions. After that, QMIX [9] utilizes a mixing network architecture to approximate all agents’ joint state-action value. The mixing network non-linearly combines each agent’s state-action value, which is only a condition on each agent’s local observations. Furthermore, QMIX enforces the monotonic joint state-action value with each agent state-action value by employing a network structural monotonicity constraint to the mixing network.
However, these algorithms cannot handle continuous action space problems. The deep deterministic policy gradient (DDPG) [6] is a representative method in the continuous action task for a single agent. Multi-agent deep deterministic policy gradient (MADDPG) [33] extended DDPG from single agent setting to the multi-agent setting by using centralized training and decentralized execution paradigm [34]. MADDPG is a representative algorithm for multi-agent continuous action problems. Thus, in this study, we model the TAPF in the intelligent warehouse as a cooperative multi-agent deep RL problem and use MADDPG to solve this problem simultaneously.
III Preliminaries
III-A Markov Decision Process and RL
This study considers a finite-horizon Markov decision process [28], defined by a tuple $(\mathcal{S},\mathcal{A},\mathcal{P},r,\gamma,T)$. $\mathcal{S}$ denotes the state space, $\mathcal{A}$ represents the finite action space, $\mathcal{P}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow[0,1]$ denotes the state transition distribution, $r:\mathcal{S}\times\mathcal{A}\rightarrow R$ denotes the reward function, $\gamma\in[0,1)$ denotes the discount factor and $T$ is a time horizon. At each time step $t$, an action $a_{t}\in A$ is chosen from a policy $\pi$. After transiting into the next state by sampling from $\mathcal{P}\left(s_{t+1}\mid s_{t},a_{t}\right)$, the agent obtains a immediate reward $r\left(s_{t},a_{t}\right)$. The agent continues performing actions until it enters a terminal state or $t$ reaches the time horizon $T$. RL aims to learn the policy $\pi:S\times A\rightarrow[0,1]$ for decision-making problems by maximizing discounted cumulative rewards $\mathbb{E}_{\pi,\mathcal{P}}\left[r_{0:T-1}\right]=\mathbb{E}_{\pi,\mathcal{P}}\left[\sum_{0}^{T-1}\gamma^{t}r\left(s_{t},a_{t}\right)\right]$. In this study, we cannot access the environment dynamics $\mathcal{P}$, which means model-free deep RL learning.
III-B Cooperative Multi-Agent Deep RL
A fully cooperative multi-agent problem can be modeled as a Markov decision process [28], described by a tuple $G=(\mathcal{S},\mathcal{A},\mathcal{P},r,\mathcal{O},N,\gamma,T)$. $s\in S$ represents the true state. At each timestep $t$, each agent $n\in N:=\{1,\ldots,N\}$ selects an action $a^{n}_{t}\in\mathcal{A}$, then the joint action $\bm{a_{t}}=\{a^{1}_{t},a^{2}_{t},...,a^{N}_{t}\}$ is obtained. $\mathcal{P}\left(s^{\prime}_{t+1}\mid s_{t},\bm{a_{t}}\right):\mathcal{S}\times\mathcal{A}^{N}\times S\rightarrow[0,1]$ denotes the state transition function. $r(s_{t},\bm{a_{t}}):\mathcal{S}\times\mathcal{A}^{N}\rightarrow R$ denotes the reward function shared by all agents, $\gamma\in[0,1)$ represents the discount factor, and $T$ denotes the time horizon. Each agent $n$ has an independent observation $o\in\mathcal{O}$ and a history of action observation $\tau^{n}\in\mathcal{T}\equiv(\mathcal{O}\times\mathcal{A})$. A stochastic policy $\pi^{n}\left(a^{n}\mid\tau^{n}\right):\mathcal{T}\times A\rightarrow[0,1]$ is based on the history of action-observation. Let $\pi^{n}$ denote agent $n$’ policy; cooperative agents aim to maximize:
$J(\pi)=\mathbb{E}_{s_{0:T-1},a^{1}\sim\pi^{1},\ldots,a^{N}\sim\pi^{N}}\left[\sum_{t=0}^{T-1}\gamma^{t}r\left(s_{t},a_{t}^{1},\ldots,a_{t}^{N}\right)\right]$ (1)
In cooperative multi-agent deep RL, the training process is centralized, but the execution process is decentralized [34]. This means that, in the learning process, each agent has access to the global state $s$. However, each agent’s decision only depends on its own action-observation history $\tau^{n}$ in the execution process.
Figure 1: Modeling TAPF as a MARL problem
Figure 2: The physical dynamics of agents
III-C Definitions of MAPF and TAPF Problems
MAPF problem: In the multi-agent path finding (MAPF) problem, the input includes an environment consisting of N agents. The environment defines all possible start and goal positions for the agents. Each agent $n$ has its own start position $start^{n}$ and goal position $goal^{n}$. At time $t$, agent $n$ is at a specific location and can execute an action $a^{n}_{t}$, subsequently moving to a new position or staying at the current position. The action can be represented as $a^{n}_{t}:\text{current position}\rightarrow\text{new position}$. The path $\pi_{n}$ of agent $n$ consists of a sequence of actions $\pi_{n}=(a^{n}_{1},a^{n}_{2},\ldots,a^{n}_{T-1})$. Conflicts in MAPF typically involve position and movement collisions between agents.
TAPF problem: The TAPF problem also involves $N$ agents within a given environment. Unlike MAPF, each agent’s goal position is not predefined. The task set includes $m$ target positions. Each agent $n$ has a defined start position, but its goal positions are dynamically assigned based on the task set. The objective of TAPF is to define a goal position assignment scheme for each intelligent system to minimize the sum of the paths of all agents. Time is also considered discrete in this problem, with agents deciding their actions at each timestep based on their current position and assigned goals.
IV Target Assignment and Path Finding for Intelligent Warehouse: A Cooperative Multi-Agent Deep RL Perspective
In Section IV-A, we detailedly describe our method that models the TAPF problem as a cooperative multi-agent RL problem. In Section IV-B, we introduce the MADDPG algorithm to solve TAPF.
IV-A Modeling TAPF as a MARL Problem
As shown in Fig. 1, we provide two scenarios to describe how we model the TAPF problem (described in Section III-C) as a cooperative MARL problem. As introduced in Section III-B, cooperative multi-agent deep RL can be modeled as a tuple $G=(\mathcal{S},\mathcal{A},\mathcal{P},r,\mathcal{O},N,\gamma,T)$. In this study, the state transition function $\mathcal{P}\left(s^{\prime}_{t+1}\mid s_{t},\bm{a_{t}}\right):\mathcal{S}\times\mathcal{A}^{N}\times\mathcal{S}\rightarrow[0,1]$ is unknown, which means the agent cannot get the environment dynamics. This is consistent with the real-world TAPF problem. The number of agents $N$ in this study can be set to any integer number. The discount factor $\gamma=0.99$ is the most common setting in deep RL. Since the main elements in multi-agent deep RL are the observation space $\mathcal{O}$ (or state space $\mathcal{S}$), the action space $\mathcal{A}$, and the reward function $r$. Thus, we describe these elements in detail.
Observation space: An agent’s observation contains the position and velocity itself, the relative positions of all tasks, the relative positions of the other agents, and the relative positions of neighboring obstacles. In this study, the number of other visible agents in an agent’s observation can be set to equal to or less than $N-1$. Taking Fig. 1 (a) as an example, the five purple circles represent five agents, and the five black circles represent five tasks. For obstacles, we set the agent to perceive the relative positions of the adjacent obstacles, and the number of adjacent obstacles can be set to any integer. In Fig. 1 (a) and (b), different color curves denote the navigation trajectories of different agents.
Figure 3: Action space of an agent
The physical dynamics of agents: Before describing the action space, we first describe the dynamics of agents. Note that our work is the first to consider the dynamics of agents in TAPF for the intelligent warehouse. As shown in Fig. 2 (a), the "No dynamics" schematic diagram represents previous literature [29, 35] methods that only focus on the next position of an agent but do not consider velocities and accelerations of agents. Further, the dynamics of agents in previous literature are solved by using traditional methods [36]. On the contrary, the "Dynamics" schematic diagram in Fig. 2 (b) represents our method that considers the dynamics of agents. We can calculate agents’ velocities and accelerations.
The output of our policy network is the magnitude and direction of four cardinal directions force ($\vec{F}_{x},\vec{F}_{-x},\vec{F}_{y},\vec{F}_{-y}$) applied to the agent. According to Newton’s second law of motion, $\vec{F}_{direction}=m*\vec{a}_{direction}$ (where $m$ is the mass of the agent), we can get the acceleration $\vec{a}_{direction}=\frac{\vec{F}_{direction}}{m}$. According to basic physics knowledge, we can get the velocity of the agent $\vec{v}_{t(direction)}=\vec{v}_{0(direction)}+\vec{a}_{direction}*\Delta t$, where $v_{0}$ is the initial velocity, and $\Delta t$ denotes a time interval.
Action space: In this study, the action space of an agent is continuous, representing the movements. As shown in Fig. 3, an agent obtains a velocity between 0.0m/s and 1.0m/s in each of the four cardinal directions [move_left ($\vec{v}_{-x}$), move_right ($\vec{v}_{x}$), move_down ($\vec{v}_{-y}$), move_up ($\vec{v}_{y}$)], the final action ($\vec{v}$) is calculated by the vector sum of the four velocities.
Reward function: The reward function is defined as:
$r=reward\_success+reward\_distance\_tasks\_to\_agents$ (2)
$\quad+reward\_collision\_agents\_to\_obstacles$
$\quad+reward\_collision\_agents\_to\_agents$
Figure 4: Distansce matrix of tasks and agents: $\|\bm{P_{tasks}-P_{agents}}\|$. The element $dis(a\ i,t\ i)$ represents the distance between $agent\ i$ and $task\ i$.
The detailed rewards are defined as follows ($R$ is the radius of an agent):
  • reward_success = $100*n$,
    wherendenotes the number of elements less than 0.05 in
[…]
  • reward_distance_tasks_to_agents
    = -$\|\bm{P_{tasks}-P_{agents}}\|_{min}$
  • reward_collision_agents_to_obstacles = -$2*n$,
    wherendenotes the number of elements less than R in
[…]
  • reward_collision_agents_to_agents = -$2*n$,
    wherendenotes the number of elements less than 2R in
[…]
For reward_success, $\|\bm{P_{tasks}-P_{agents}}\|$ represents the distance matrix of all tasks to all agents. Fig. 4 shows a detailed distance matrix. If the distance of $agent\ i$ to $task\ i$ is less than $0.05m$, we give a $+100$ positive reward to force agents to navigate to the tasks. If there are $n$ elements in $\|\bm{P_{tasks}-P_{agents}}\|$ less than $0.05m$, we set reward_success = +$100*n$. Reward_distance_tasks_to_agents is created to provide a dense reward to accelerate the training speed. In $\|\bm{P_{tasks}-P_{agents}}\|_{min}$, the subscript $min$ means that we sum the minimum distance of each task to every agent.
Reward_collision_agents_to_obstacles aims to punish agents from collisions with obstacles. The definition of $\|\bm{P_{obstacles}-P_{agents}}\|$ is similar with $\|\bm{P_{tasks}-P_{agents}}\|$, $\|\bm{P_{obstacles}-P_{agents}}\|$ denotes the distance matrix of obstacles and tasks. If the distance of an agent to an obstacle is less than $R$, a negative reward $-2$ is given. If there are $n$ elements in $\|\bm{P_{obstacles}-P_{agents}}\|$ less than $R$, we set reward_collision_agents_to_obstacles = -$2*n$.
Reward_collision_agents_to_agents aims to punish agents from collision to agents. The definition of $\|\bm{P_{agents}-P_{agents}}\|$ is similar with $\|\bm{P_{tasks}-P_{agents}}\|$, $\|\bm{P_{agents}-P_{agents}}\|$ denotes the distance matrix of agents and agents. If the distance of an agent to an agent is less than $2R$, we give a negative reward $-2$. If there are $n$ elements in $\|\bm{P_{agents}-P_{agents}}\|$ less than $2R$, we set reward_collision_agents_to_agents = -$2*n$.
IV-B Solving TAPF Problem via Cooperative MARL
In this study, we introduce MADDPG [33] to solve the TAPF problem. Considering that agents are homogeneous, they can share the same policy network. This makes learning policy more efficient. Let $N$ denote the number of agents, and $\pi$ denote the policy parameterized by $\theta$. The centralized critic $Q$ is parameterized by $\phi$. The parameters $\theta$ for policy $\pi$ can be updated iteratively by the associated critic $Q$:
$\nabla_{\theta}J\left(\theta\right)$ $=\mathbb{E}_{\left(\bm{x_{t}},\bm{a_{t}}\right)\sim\mathcal{D}}\Big{[}\nabla_{\theta}\pi\left(o_{t}^{n}\mid\theta\right)$ (3)
$\quad\nabla_{a_{t}}Q\left(\bm{x_{t}},a_{t}^{1},\ldots,a_{t}^{N}\mid\phi\right)|_{a_{t}^{n}=\pi\left(o_{t}^{n}\mid\theta\right)}\Big{]}$
where $\bm{x}_{t}$ and $\bm{a}_{t}$ are the concatenation of all agents’ observation and action at time step $t$, $\bm{x}_{t}=\left(o_{t}^{1},\ldots,o_{t}^{N}\right)$ and $\bm{a}_{t}=\left(a_{t}^{1},\ldots,a_{t}^{N}\right)$. $o_{t}^{n}$ is the observation received by agent $n$ at time step $t$. $\mathcal{D}$ denotes the replay buffer containing tuples ($\bm{x}_{t},\bm{a}_{t},r_{t}\cdots$).
The centralized critic parameters $\phi$ are optimized by minimizing the loss:
$L\left(\phi\right)=\mathbb{E}_{\left(\bm{x}_{t},\bm{a}_{t},{r}_{t},\bm{x}_{t+1}\right)\sim\mathcal{D}}\left[\left(Q\left(\bm{x}_{t},\bm{a}_{t}\mid\phi\right)-y_{t}\right)^{2}\right]$ (4)
where ${r}_{t}$ is the concatenation of rewards (Eq (2)) received by all agents at time step $t$. The target critic value $y_{t}^{n}$ is defined as:
$y_{t}=r_{t}+\gamma Q^{-}\left(\bm{x}_{t+1},a_{t+1}^{1},\ldots,a_{t+1}^{N}\mid\phi^{-}\right)|_{a_{t+1}^{n}=\pi\left(o_{t+1}\right),n=1,\ldots,N}$ (5)
where $Q^{-}$ is a target Q-network parameterized by $\phi^{-}$. $\phi^{-}$ is optimized by:
$\phi^{-}=\alpha\phi+(1-\alpha)\phi^{-}$ (6)
where $\alpha$ is a coefficient to trade off the weight of Q-network and target Q-network.
The shared reward makes the agents learn policy cooperatively. In the execution phase, we only use the policy $\pi$ whose inputs are $o_{t}^{1},\ldots,o_{t}^{N}$, respectively, and the outputs are force ($\bm{\vec{F}}$) applied to the agent. According to the physical dynamics of agents described in Section IV-A, we can get the actions $\bm{a}_{t}=\left(a_{t}^{1},\ldots,a_{t}^{N}\right)$. Note that, for agent $n$, the input of $\pi$ is only the observation $o_{t}^{n}$ received by agent $n$ at time step $t$ in the execution phase.
Figure 5: Average return
V Experiments
In this section, we verified the proposed method from three aspects. In Section V-A, we verified the proposed method’s target assignment and path planning performances. In all the experiments, the locations of agents and task points are generated randomly. Section V-B verified the learned cooperation ability. Section V-C verified the time efficiency.
Figure 6: Two agents - two tasks
V-A Target Assignment and Path Planning
This subsection verified the proposed method in various intelligent warehouse settings. We set five different level scenarios to verify the performances: (1) two agents - two tasks (2) two agents - four tasks (3) five agents - five tasks (4) five agents - ten tasks (5) five agents - twenty tasks.
Fig. 5 shows the training average return curves. In all the different level scenarios, the average returns increase monotonously. This verifies the stability of the proposed method. Fig. 6 to Fig. 10 show the performance target assignment and pathfinding in five different level scenarios. Experimental results show that although the difficulty of these five different level scenarios increases gradually, the proposed method performs well in all the tasks.
Figure 7: Two agents - four tasks
For easy tasks (two agents - two tasks) in Fig. 6 (a) (b) (c) (d), we can see that the target assignment and pathfinding were addressed well. For target assignment, results show that the task assignment is addressed very reasonably because the proposed method assigns the task to the agent that is close to the task. For pathfinding, it is shown that the planned paths are almost shortest.
Figure 8: Five agents - five tasks
For difficult tasks (five agents - twenty tasks) in Fig. 10 (a) (b) (c) (d), we can see that the target assignment and pathfinding were also addressed well. For target assignment, it can be seen that each agent is reasonably assigned to several nearby tasks. For pathfinding, it is shown that the planned paths are usually the shortest. The same superior performances can be seen in other different level scenarios, such as two agents - four tasks in Fig. 7, five agents - five tasks in Fig. 8, and five agents - ten tasks in Fig. 9. This demonstrated the efficiency of the proposed method.
Figure 9: Five agents - ten tasks
Figure 10: Five agents - twenty tasks
V-B Cooperation Ability
We also designed a conflict scenario to verify the agents’ learned cooperation ability. As shown in Fig. 11, the big purple circle represents agent-1, its task is denoted by the small purple circle. The big gray circle represents agent-2, and the small black circle denotes its task. We deliberately block other roads to create a conflict environment. As shown in Fig. 11, there is bound to be a conflict between agent-1 and agent-2 during the navigation. The trajectories generated by the proposed method are shown in Fig. 11. The red curve is the trajectory of agent-1, and the cyan curve is the trajectory of agent-2. Results show that both agent-1 and agent-2 learned to avoid each other at the point of conflict and then navigated to their tasks. This verified the cooperation ability of the proposed method.
TABLE I: Time comparisons between our method and traditional methods
Scenarios Ours Traditional methods
two agents - two tasks 0.0003221s(TAPF-ES) 0.02091s(TA)+0.0038107s(PF-ES)
two agents - four tasks 0.0003287s(TAPF-ES) 0.03887s(TA)+0.0036540s(PF-ES)
five agents - five tasks 0.0003375s(TAPF-ES) 0.0498669s(TA)+0.0037150s(PF-ES)
five agents - ten tasks 0.0003414s(TAPF-ES) 0.2918257s(TA)+0.0054604s(PF-ES)
five agents - twenty tasks 0.0003505s(TAPF-ES) 14.66s(TA)+0.0038555s(PF-ES)
Figure 11: The cooperation of agents
V-C Time Efficiency
Table I shows the time comparisons between traditional methods [37, 38] and the proposed method. Table I shows that the proposed method simultaneously addresses target assignment (TA) and path planning (PF) problems. However, the traditional method addresses target assignment first and then performs path planning. Thus, the time consumed by the proposed method is TAPF time. However, the time consumed by the traditional method is TA+PF. In Table I, ES means each step. The reasons that we compare the consumed time of each step are: (1) Our method handles continuous action space problems. However, the traditional method handles discrete action space problems. (2) The consumed time of path finding can be influenced by many factors, such as the resolution of the grid map and the distance taken at each step in the grid map. (3) Previous literature [37, 38] did not consider the physical dynamics of agents. However, the physical dynamics of agents have been considered in this study. Thus, for a fair comparison, we only compare the consumed time of each step, which means the consumed time from an agent obtains an observation to output the decision action. For the traditional method, the consumed time of target assignment is computed by the method in [37], and the consumed time of each step in path planning is computed by the method in [38].
Table I shows that in easy tasks such two agents - two tasks and two agents - four tasks, the consumed time of target assignment in the traditional method can be acceptable. However, as the difficulty of tasks increased, the time of solving target assignments in the traditional method increased rapidly, especially in the five agents - twenty tasks scenario. However, the proposed method provides policy time-efficiently in all different level tasks (from two agents - two tasks to five agents - twenty tasks). This verified the time efficiency ability of the proposed method. As we all know, time efficiency is an important factor in real-world TAPF problems. Therefore, the experiment results demonstrated the promising engineering practical value of the proposed method. Note that Fig. 6 to Fig. 10 had shown the performances of target assignment and path finding of the proposed method in five different level scenarios. For target assignment, results show that the task assignment is addressed reasonably because the proposed method assigns the task to the agent that is close to the task. For pathfinding, results show that the planned paths are almost shortest.
VI Conclusion and future work
This paper proposed a method to simultaneously solve the TAPF problem in the intelligent warehouse from a cooperative multi-agent deep RL perspective. First, we modeled the TAPF problem as a cooperative multi-agent deep RL problem. Then, we simultaneously solved task assignment and path finding using a cooperative multi-agent deep RL algorithm. Moreover, previous literature rarely considers the physical dynamics of agents. In this study, the physical dynamics of agents have been considered. Experimental results showed that the proposed method performed well in various task settings, which meant the target assignment was solved reasonably, and the planned path was almost the shortest. Furthermore, the proposed method is more time-efficient than baselines. For future work, we will apply the proposed method to real-world TAPF problems.